# Repeating Flares, X-ray Outbursts and Delayed Infrared Emission: A Comprehensive Compilation of Optical Tidal Disruption Events

### v1

@@ 4.3 Repeating flares
The bottom panels of Fig. 6 illustrate the multiple flares of each TDE, overlaid for direct comparison. We normalize the flux to the corresponding flare peak for comparison. We observe that the TDEs in the repeating flare subsample display flares that are similar in shape with the confirmed TDE flares.
In the following, we examine in detail the repeating flare TDEs that have not previously been studied extensively, namely: AT 2024pvu, AT 2022exr, AT 2021uvz, AT 2020acka, AT 2019teq and AT 2019ehz. The remaining repeating flare TDEs are discussed in Appendix C.
4.3.1 AT 2024pvu
AT 2024pvu showed a previous outburst approximately 17.9 years prior to the 2024 flare. The first flare was detected by the CRTS, while the second, most recent flare was observed by the ZTF survey in all three available bands (, and ). The two flares are plotted in the upper panel of Fig. 6(a), while their shapes are compared in the bottom panel. A visual inspection, reveals that the two flares exhibit similar shapes.
Unfortunately, due to limited photometric and spectroscopic data of the initial flare, we cannot confirm whether it was caused by a TDE or another type of transient. Furthermore, there are no spectra available for the host galaxy, or any additional information indicating that it is a post-starburst galaxy, where as we discuss later, TDE rates are significantly enhanced. However, we can confidently exclude the possibility that both flares resulted from the disruption of the same star. If we assume a solar-like star, a SMBH with a mass of and an orbital period of 17.9 years, an orbital eccentricity of 0.99932 would be required. Such an orbit would be highly unstable, making this scenario unlikely.
4.3.2 AT 2022exr
AT 2022exr displays one of the most unique light curves in our sample, with a distinct double peak, followed by a third, smaller peak as (see in the upper panel of Fig. 6(b)). The ZTF -band light curve highlights in detail all three flares, even though there is a gap in coverage during the rise of the first flare. All three flares occurred within 250 days, with the second flare appearing before the first one had fully faded and the third flare before the second had ended.
Since no publicly available spectra exists for the second and third flares, we cannot confirm their TDE nature. However, likely all three flares are related to TDE activity. All three normalized light curves exhibit a similar shape, albeit with different peak fluxes. As we discuss in more detail in Sect. 5, this feature is characteristic of repeating TDE related flares.
We assume that all three flares are TDE-related, as it is highly unlikely that they are caused by separate transient events. We can study the partially-disrupted TDE (pTDE) and double TDE (disruption of a stellar binary) scenarios, since they are considered singular events. The pTDE case is unlikely due to inconsistent timing of the flare intervals: the first and second flares occur roughly 110 days apart, while the second and third are separated by about 35 days. Such a dramatic change in timing would require the star’s orbit to be significantly altered. This inconsistency could be explained by a pTDE for the first two flares and accretion disc emission for the third flare.
An alternative interpretation for the observed light curve morphology is the double TDE scenario, followed by a pTDE. In double TDEs, a stellar binary is disrupted by the SMBH (e.g. Mandel & Levin 2015; Mainetti et al. 2016). Mandel & Levin (2015) note that double TDEs could produce a distinct double-peaked signature, where one star is initially disrupted, while the second is captured in an elliptical orbit around the SMBH. This scenario could potentially explain for AT 2022exr, with the first star being fully disrupted and the second one entering an elliptical orbit undergoing partial disruptions, producing the second and third flares. However, Mandel & Levin (2015) found through simulations that the orbital period of the second star can range from 6 months up to several decades, with a median period of roughly 50 years. In the case of AT 2022exr the time interval between the first two flares is only 110 days, shorter than the expected lower limit, making this scenario less likely. Furthermore, Mainetti et al. (2016) modelled light curves of double TDEs using smoothed particle hydrodynamics simulations; however, these simulations did not produce double peaked light curves. This further challenges the double TDE interpretation for AT 2022exr.
4.3.3 AT 2021uvz
Similar to AT 2022exr, AT 2021uvz displays a double-peaked flare, as seen in the upper panel of Fig. 6(e). We observe that both peaks share the same shape, despite their close temporal proximity.
Following our discussion of AT 2022exr’s double-peaked flare, it is highly unlikely that two separate events caused AT 2021uvz’s flares. Under the repeating pTDE hypothesis, we can assume an orbital period of roughly 55 days (the time interval between peaks), which is theoretically possible, with the star being completely disrupted after its second arrival at the pericenter. As for the double TDE scenario, the light curve of AT 2021uvz more closely resembles what one would expect if both stars in the binary system were tidally disrupted almost immediately one after the other. Mandel & Levin (2015) found that 20% of their simulated double TDEs had a time difference in peak luminosities times exceeding 2 months, with individual rise times of less than 1 month. However, the first flare has an observed rise time of 50 days. We note that Mandel & Levin (2015) define the rise time as the duration required for the accretion rate to increase from low values to its peak for each star’s disruption, whereas we refer to the time taken for the light curve to rise from the non-outburst plateau to peak luminosity. Since the origin of optical emission in optical TDEs is not clear, we cannot further assess the plausibility of this scenario.
4.3.4 AT 2020acka
The upper panel of Fig. 7(a) displays the ZTF -band light curve of the AT 2020acka’s flares. The second flare appears more like a bump than a flare, which could result from an accretion disc fluctuation or a sudden influx of excess orbiting material from the disrupted star. The nature of this outburst is not clear from the -band light curve alone. To investigate further, we studied the ZTF -band light curve shown in Fig. 8. Despite the large gap in the -band data, the second flare is more prominent. From the bottom panel of Fig. 7(a) and Fig. 8 we cannot conclusively determine whether the two flares share similar shapes. Consequently, we cannot reliably comment on the nature of the second, bump-like flare.
4.3.5 AT 2019teq
Approximately 300 days prior to the TDE flare, this source exhibited unexpected variability. Although it was included in the Hammerstein et al. (2023) sample, the pre-TDE flare-like feature was not addressed. We present both the ZTF -band and -band light curves in the upper panels of Figs. 7(c) and 7(d) respectively, as artifacts can sometimes affect only one band. Unfortunately there is a gap in the data after the rise of the first flare. However it is evident that the rise time differs significantly from the TDE flare resembling the behaviour observed in AT 2021mhg.
Interestingly, AT 2019teq was detected in X-rays both during the TDE flare period and 3 years later (Yao & Guolo 2022; Ajay et al. 2022). It was also detected in the radio at 6 GHz using the Very Large Array (VLA) (Cendes et al. 2022). Yao & Guolo (2022) and Ajay et al. (2022) suggest that the observed brightening and spectral hardening of the X-rays indicate the formation of a corona. The radio detection further suggests the possible launch of a jet, as only upper limits were obtained in two previous observations.
4.3.6 AT 2019ehz
This TDE was first studied as a part of the van Velzen et al. (2021) sample and was later included in both the Hammerstein et al. (2023) and Yao et al. (2023) samples. However, none of these studies addressed its second flare-like structure. The light curves for both the ZTF -band and -band are presented in the upper panels of Figs. 7(c) and 7(d) respectively. In the bottom panels of these figures, we compare the shapes of the two flares in each filter. Notably, in the -band light curve both flares exhibit a similar shape, consistent with the other repeating TDE-related flares in our sample. However, this is not the case for the -band light curve, where the shapes differ. This behaviour is reminiscent of the flares of AT 2020acka, where the second flare-like bump was more prominent in the -band in than the -band. This behaviour could be due to accretion disc emission but the similarity of the flare shapes in the -band point toward an alternative explanation.
Under the assumption that the second flare is related to the initial TDE, we can investigate its repeating nature. We first assess the likelihood that the two flares resulted from two separate stars being disrupted in the same galaxy. AT 2019ehz was one of the 33 TDEs in the sample of Yao et al. (2023), from which they estimated an optical TDE rate of yr-1 galaxy-1. Following the methodology of Sun et al. (2024), we estimate the probability of a TDE occurring after the initial flare as . Adopting the aforementioned optical TDE rate and the time interval between flares to be days, we obtain . This probability is extremely low. For to exceed , the TDE rate would need to be roughly 1900 times higher, which is far beyond observed values, even for the theoretically enhanced rates in post starburst galaxies. Hence we can confidently exclude the scenario that the two flares were caused by separate events.
In the pTDE case the stellar remnant would need to be in an orbit with a period of 300 days, which is theoretically possible. However, the time separation between the two flares is too large for them to be caused by a full double TDE. The only viable scenario would involve the first star being fully disrupted while the second star is caught in an elliptical orbit around the SMBH, with an orbital period similar to that required for the repeating pTDE interpretation. We note that the first flare’s light curve in Fig. 7(e) displays a knee141414Noticeable change in the slope of the light curve. in its decay, resembling features found in the simulations of Mainetti et al. (2016). In these simulations, such features were only seen in cases where both stars in an equal-mass binary were partially disrupted or in cases with highly unequal-mass binaries.
4.4 Infrared Flares
A small subset (18) of TDEs in our sample display IR flares, along with their optical counterpart, namely : AT 2023ugy, AT 2023cvb, AT 2022upj, AT 2022fpx, AT 2022agi, AT 2022dyt, AT 2022aee, AT 2021uqv, AT 2020afhd, AT 2020ksf, AT 2020mot, AT 2020nov, AT 2020pj, AT 2019qiz, AT 2019dsg, AT 2019azh, AT 2017gge and ASASSN-14li.
@@ 4.4 Infrared Flares
In Fig. 9 we show two examples of infrared light curves (lower panels) from the NEOWISE survey (W1 and W2 filters), in addition to optical (middle panels) and X-ray data (upper panels). Only 4 TDEs, namely AT 2023ugy, AT 2022aee, AT 2020mot, AT 2020pj, lacked an X-ray counterpart (with only upper limits available). The rest of the light curves can be plotted using the accompanying app that is available on the GitHub page.
4.5 UV Emission
@@
Following the process described in Sect. 3.3, we retrieved UVOT photometry for 118 TDEs. An example light curve (AT 2020nov) is displayed in Fig. 10. In total, 111 TDEs have data for all 6 wavebands, while 95 have at least 5 observations for each of the three UV bands. These data can be accessed through the GitHub page, under the UVOT folder in the photometry section. Additionally, these light curves can be plotted using the accompanying app, also available on the GitHub page.
4.6 X-ray outbursts
@@
We also investigated TDEs observed in X-rays (see Sect. 3.2). In our main sample, 122 sources have either detections or 3 upper limits in the X-rays. Of these, 45 TDEs have at least 1 detection, while 26 have at least 5 detections. Example light curves are shown in the upper panels of Fig. 9. All available X-ray data can be found in the accompanying GitHub repository and plotted using the provided app. We note that we analyse X-ray observations individually, meaning that stacking observations will lead to more detections. We also consider as detections only the observations with signal-to-noise (SNR) ratio equal to or greater than 3. The SNR of the observations is also included in the GitHub files of each source, so the users can set the limit to the threshold they want.
@@ 5.1 Statistics of the sample
In Sect. 4.1, we showed that the distributions of the duration, and , as well as their ratio, are consistent with a log-normal distribution. Log-normal distributions are usually encountered in systems driven by multiplicative processes or exponentially growing or decaying dynamics. They also emerge when the observed data result from a combination of independent processes with varying initial conditions. This means that the diversity in stellar properties and BH environments in TDEs could be responsible for the log-normal distributions.
The log-normal behaviour of the redshift distribution could result from observational selection effects rather than from some intrinsic property of TDEs. We expect that the distribution of redshifts will change with the addition of data from upcoming surveys such as the Legacy Survey of Space and Time151515https://www.lsst.org/about (LSST; Ivezic et al. 2008).
@@ 5.2 Spectral classes
As shown in Fig. 5, the majority of TDEs fall into the TDE-H+He category. As predicted by Syer & Ulmer (1999), the disruption of main-sequence (MS) stars by SMBHs of mass would be the easiest for modern-day telescopes to observe. MS stars consist of both hydrogen and helium, hence the observed spectral signatures. Moreover, MS stars are abundant in galactic centres due to their long-lived nature and have been found to dominate TDE rates in past studies (e.g. Kochanek 2016). The second most common spectral class in our sample is TDE-H, followed by TDE-He and TDE-featureless classes. The statistical study of Nicholl et al. (2022) uncovered that for the TDE-H sources in their sample, their impact parameter (b; i.e. high b almost full disruption, small b partial disruption) was by far the smallest compared to TDE-H+He and TDE-He, corresponding to partial disruptions. The outer layers of MS stars are mainly hydrogen, so the partial disruption of the outer layers of MS stars could explain the observed TDE-H population. Additionally, hydrogen-rich young MS stars are typically rare, hence the contribution of their disruption to the population should be small. On the other hand, the TDE-He class is significantly more scarcely observed, with only 16 events (18 if we consider the different epoch classifications from Sect. 4.2). This could be explained by the disruption of helium-rich stellar cores, as suggested by Gezari et al. (2012) for PS1-10jh. These stellar remnants are rare, theorized to be originating from stellar binary systems (e.g. Paczyński 1967; Iben & Tutukov 1985; Götberg et al. 2019), which could explain their small observed number in TDEs.
@@
Regarding the TDE-featureless class, we have collected data for 13 such events. Hammerstein et al. (2023) found that the 4 TDE-featureless sources in their sample exhibited higher peak bolometric luminosities, hotter temperatures and larger radii than those in the TDE-H and TDE-H+He classes. Moreover, their TDE-featureless sample favoured redder and more massive host galaxies compared to other classes. This trend can be further tested using the TDEs in our catalogue. Featureless TDE spectra can arise for several reasons. One possibility is that the line-emitting regions are obscured, depending on our viewing angle. Another explanation is that dense, optically thick outflows reprocess and thermalize high-energy photons, effectively ”washing out” sharp emission lines. Additionally, a strong, hot continuum can overpower weaker lines, making them difficult to detect.
@@
To analyse the variability properties of different spectral classes, we applied the same statistical approach used in Sect. 4.1 to the TDEs for which we could reliably measure flare variability. We then utilized a K-S test to assess whether the distributions of different spectral classes are consistent with that of the full sample. We obtained p-values ranging from 0.15 to 0.97 for the rise times, decay times and durations. Although the sample sizes for TDE-H, TDE-He and TDE-featureless classes are small, we find that the time scale distributions of all spectral classes are drawn from the same parent population.
@@
Additionally, we investigated the spectral classes of TDEs in different subsamples showing either IR emission, repeating flares or X-ray emission. For the IR sample we find that 12 are TDE-H+He, 5 are TDE-H and 1 is TDE-He. For the repeating flare sample we find that 8 are TDE-H+He, 2 are TDE-H and 2 are TDE-featureless. Finally, for the X-ray detection sample we find that 4 are TDE-He, 26 are TDE-H+He, 9 are TDE-H and 3 are TDE-featureless. We estimated the probability that the spectral classes of the different subsamples are randomly drawn from the full sample using:
@@
where is the binomial coefficient. Here, the numerator represents the number of ways to compose a sample with the different spectral classes in each subsample while the denominator accounts for all possible selections from the full set of 134 TDEs. The small p-values in all three cases indicate the spectral composition of these subsamples is unlikely to be random. This could indicate that the spectral characteristics of TDEs are related to their overall multi-wavelength behaviour.

### v2

@@ 4.3 Repeating flares
The bottom panels of Fig. 6 illustrate the multiple flares of each TDE, overlaid for direct comparison. We normalise the flux to the corresponding flare peak for comparison. We observe that the TDEs in the repeating flare sub-sample display flares that are similar in shape with the confirmed TDE flares. Most notably, we find AT 2024pvu to be the record holder for the longest time separation between two flares ( yr), as well as that AT 2022exr and AT 2021uvz display a double-peak morphology.
For each of the repeating flare TDEs, we examine the possible mechanisms that could explain their multiple flares. We investigate the repeating partial TDE scenario (pTDE), the disruption of a stellar binary (double TDE) and the disruption of separate, unrelated stars. We also discuss alternative mechanisms for specific cases. A summary of this analysis is presented in Table 2, where the first three columns show the name, the time separation between flares (; taken from peak to peak) and the number of flares () respectively, while the final four columns display the different scenarios we considered for the repeating nature of the events (likely, could explain the event but not very probable, not likely). The relevant references are listed in the final column.
First we explored the separate event scenario for the 11 sources in the repeating flare sample. In order to achieve this, we follow the analysis performed in Sun et al. (2024) for AT 2022agi (IRAS F01004-2237; see their Sect. 5.4). The probability of a separate event occuring in the same host after a time interval of can be computed by , where yr-1 galaxy-1, which is the incident optical TDE rate from Yao et al. (2023). For each TDE we consider the displayed in the second column of Table 2, whereas the probability for each host galaxy to have another separate event in is given in the parenthesis of the sixth column. We observed that the smallest value is , while the largest is . For to exceed , the TDE rate would need to be roughly 87, 152 and 108 times higher for AT 2024pvu, AT 2022agi and AT 2019azh respectively, while for the rest it would need to be 422 times higher or more, reaching times for AT 2021uvz. For the first three we list, the required rate could be theoretically possible given the enhanced rates in post starburst galaxies, which can have boosted rates of up to times on average (e.g. French et al. 2016; Law-Smith et al. 2017; Graur et al. 2018). However, for the latter cases, we can confidently exclude the scenario that the two flares were caused by separate events, since their rates would need to be enhanced by more than 400 times. We note that the reported p-values do not account for the look-elsewhere effect and could be substantially higher (even by a factor of 100 depending on the event). The correct treatment of the look-elsewhere effect would require dedicated simulations which are beyond the scope of this work.
In order to study whether the multiple flares could be explained by the partial disruption of the same star, we utilised a toy model to calculate the eccentricity. If we assume a solar-like star, a SMBH with a mass of and an orbital period of , we retrieve orbits between the range and . Such orbits would be highly unstable, making the pTDE scenario unlikely. However, such highly eccentric orbits could be explained by the Hills Mechanism (Hills 1988), where the SMBH tidally breaks a stellar binary: one star is ejected with hyper-velocity while the other is captured on an extreme eccentric bound orbit. We use another toy model to qualitatively probe this scenario. Using Eq. 3 of Pfahl (2005, or Eq. 8 of ) the semi-major axis of the captured star can be associated with the properties of the binary (the binary semi-major axis , , and the total mass of the binary, ). Assuming two solar-like stars and , we find that using as the observed period yields values that are between to . We can access if the two stars are far enough to avoid a common envelope by calculating the effective radius of the Roche lobe , , using Eq. 2 in Eggleton (1983). For two cases, namely AT 2022exr and AT 2021uvz (which have a double-peaked light curve morphology), we find that this scenario is not plausible, since . To further probe the parameter space of the toy model, we use an range [] with a step of 0.5 dex, as well as unequal mass and radius stellar binaries. For each case we repeat the above analysis and we also consider only the cases where . We find that all the multiple flares could be explained by our toy model in the aforementioned parameter space, and hence we cannot exclude the pTDE scenario for any of them.
| Name | (days) | pTDE | double TDE | separate events (p) | Other | References | |
| AT 2024pvu | 2 | () | - | i | |||
| AT 2022exr | () | 3 | ? | X () | - | i | |
| AT 2022dbl | 2 | X () | - | ii, iii | |||
| AT 2022agi | 2 | () | - | iv | |||
| AT 2021uvz | 2 | ? | X () | - | i | ||
| AT 2021mhg | 2 | X | X | X () | v, vi | ||
| AT 2020vdq | 2 | X () | - | vii, viii, ix | |||
| AT 2020acka | 2 | X () | x, xi | ||||
| AT 2019azh | 2 | () | - | xii, xiii | |||
| AT 2019ehz | 2 | X () | vii, xi, xiii | ||||
| AT 2018fyk | 2 | X () | - | xiv, xv |
An alternative explanation of the repeating flares could be the disruption of a stellar binary (double TDE), which has been studied through simulations (e.g. Mandel and Levin 2015; Mainetti et al. 2016). The difference with the Hills mechanism is that now both stars are disrupted, either in sequence or with a delay, since after the first TDE, the second star is captured in an elliptical orbit around the SMBH. Mandel and Levin (2015) use a population model for the binaries approaching the SMBH with small impact parameters and through numerical experiments, which follow Newtonian stellar dynamics of binaries that approach a SMBH. They found that 18 of their simulations yield the subsequent disruption of both stars. As noted by Mandel and Levin (2015), this channel could produce a double-peaked flare similar to the ones observed in AT 2022exr and AT 2021uvz, but they stress that hydrodynamical simulations would be needed to test this hypothesis. Mainetti et al. (2016) modelled light curves of double TDEs using smoothed particle hydrodynamics simulations; however, these simulations did not produce double peaked light curves. Nevertheless they did find a knee171717Noticeable change in the slope of the light curve. in the decay of the simulated light curves in the cases where both stars in an equal-mass binary were partially disrupted or in cases with highly unequal-mass binaries. We observe a knee in the decay of the light curve of AT 2019ehz (see Figs. 7(c) and 7(d)), where one of the stars could have been caught in an orbit to explain the second flare days after the first one. Furthermore, in about of their simulations, Mandel and Levin (2015) found that one star would get fully disrupted, while the second was captured in a bound orbit with a period of at least 6 months (and a median period of roughly 50 years). This scenario could explain the double peaked light curves of all the repeating flare TDEs, except AT 2022exr, AT 2021mhg and AT 2021uvz.
There are also additional mechanisms that might be responsible for the multiple flares. One example is AT 2021mhg, where the TDE flare was followed by a SN Type Ia (Somalwar et al. 2025b). Additionally, AT 2019ehz and AT 2020acka display secondary, dimmer flares, which could result from an accretion disc fluctuation or a sudden influx of excess orbiting material from the disrupted star (Guo et al. 2025).
4.4 Multi-wavelength flares
@@
A small subset (18) of optical TDEs in our sample display IR flares, namely, AT 2023ugy, AT 2023cvb, AT 2022upj, AT 2022fpx, AT 2022agi, AT 2022dyt, AT 2022aee, AT 2021uqv, AT 2020afhd, AT 2020ksf, AT 2020mot, AT 2020nov, AT 2020pj, AT 2019qiz, AT 2019dsg, AT 2019azh, AT 2017gge, and ASASSN-14li. In Fig. 8 we show two examples of IR light curves (lower panels) from the NEOWISE survey (W1 and W2 filters), in addition to optical (middle panels) and X-ray data (upper panels). Only four TDEs (i.e. AT 2023ugy, AT 2022aee, AT 2020mot, and AT 2020pj) lack an X-ray counterpart (only upper limits are available).
@@
Following the process described in Sect. 3.3, we retrieved UVOT photometry for 118 TDEs. An example light curve (AT 2020nov) is displayed in Fig. 9. In total, 111 TDEs have data for all six wavebands, while 95 have at least five observations for each of the three UV bands. These data can be accessed through the GitHub page, under the UVOT folder in the photometry section.
@@
We also investigated TDEs observed in X-rays (see Sect. 3.2). In our main sample, 126 sources either have detections or 3 upper limits in the X-rays. Of these, 45 TDEs have at least one detection, while 26 have at least five detections. Example light curves are shown in the upper panels of Fig. 8. We note that we analyse X-ray observations individually, meaning that stacking observations will lead to more detections. We also consider as detections only the observations with signal-to-noise (S/N) ratio equal to or greater than three. The S/N of the observations is also included in the GitHub files of each source, so the users can set the limit to the threshold they want. All available X-ray, UV and IR data can be found in the accompanying GitHub repository and plotted using the provided app.
@@ 5.1 Statistics of the sample
In Sect. 4.1, we showed that the distributions of the duration, and , as well as their ratio, are consistent with a log-normal distribution. These distributions, along with the scatter plot in the bottom-right panel of Fig. 3, can be used to generate mock TDE data. We note that this is the first time that the timescales in TDEs have been parametrised with a large sample of more than 100 TDEs. Previously, van Velzen et al. (2021) had a final sample of 39 TDEs, for which they inferred their timescales through fitting, and found that there was no correlation between rise times and decay times. However, Hammerstein et al. (2023) and Yao et al. (2023), who studied 30 and 33 TDEs respectively, later reported that there is a positive correlation between rise and decay times, similar to our results (bottom-right panel of Fig. 3).
The log-normal behaviour of the redshift distribution could result from observational selection effects, due to the sensitivity of current telescopes and the local universe galaxy density, rather than from some intrinsic property of TDEs. We expect that the distribution of redshifts will change with the addition of data from upcoming surveys such as the Legacy Survey of Space and Time181818https://www.lsst.org/about (LSST; Ivezic et al. 2008). Furthermore, the redshift distribution, along with the rest of the distributions from Sect. 4.1, can be used to generate mock TDE populations and light curves.
@@ 5.2 Spectral classes
As shown in Fig. 5, the majority of TDEs fall into the TDE-H+He category. As predicted by Syer and Ulmer (1999), the disruption of main-sequence (MS) stars by SMBHs of mass would be the easiest for modern-day telescopes to observe. MS stars consist of both hydrogen and helium, hence the observed spectral signatures. Moreover, MS stars are abundant in galactic centres due to their long-lived nature and have been found to dominate TDE rates in past studies (e.g. Kochanek 2016). The second most common spectral class in our sample is TDE-H, followed by TDE-He and TDE-featureless classes. The statistical study of Nicholl et al. (2022) uncovered that for the TDE-H sources in their sample, their impact parameter (b; i.e. high b almost full disruption, small b partial disruption) was by far the smallest compared to TDE-H+He and TDE-He, corresponding to partial disruptions. The outer layers of MS stars are mainly hydrogen, so the partial disruption of the outer layers of MS stars could explain the observed TDE-H population. Additionally, hydrogen-rich young MS stars are typically rare, and hence the contribution of their disruption to the population should be small. On the other hand, the TDE-He class is significantly more scarcely observed, with only 16 events (18 if we consider the different epoch classifications from Sect. 4.2). This could be explained by the disruption of helium-rich stellar cores, as suggested by Gezari et al. (2012) for PS1-10jh. These stellar remnants are rare, theorised to be originating from stellar binary systems (e.g. Paczyński 1967; Iben and Tutukov 1985; Götberg et al. 2019), which could explain their small observed number in TDEs (e.g. Mockler et al. 2024).
@@
Regarding the TDE-featureless class, we have collected data for 13 such events. Hammerstein et al. (2023) found that the 4 TDE-featureless sources in their sample exhibited higher peak bolometric luminosities, hotter temperatures and larger radii than those in the TDE-H and TDE-H+He classes. Moreover, their TDE-featureless sample favoured redder and more massive host galaxies compared to other classes. This trend can be further tested using the TDEs in our catalogue. Additionally, the Hammerstein et al. (2023) featureless TDEs were found in relatively higher redshifts than the rest of their sample. Similarly, we find that 7 of our 13 featureless TDEs display redshifts that are 0.25 or above, where the mean is around (median around 0.09). In our sample, is at the 91.6th percentile for the optical TDEs (grey filled histogram from Fig. 4), meaning that 8.4% of the optical TDEs have . Featureless TDE spectra can arise for several reasons. One possibility is that the line-emitting regions are obscured, depending on our viewing angle. Another explanation is that dense, optically thick outflows reprocess and thermalise high-energy photons, effectively ‘washing out’ sharp emission lines. Additionally, a strong, hot continuum can overpower weaker lines, making them difficult to detect.
@@
To analyse the variability properties of different spectral classes, we applied the same statistical approach used in Sect. 4.1 to the TDEs for which we could reliably measure flare variability. We then utilised a KS test to assess whether the distributions of different spectral classes are consistent with that of the full sample. We obtained p-values ranging from 0.15 to 0.97 for the rise times, decay times and durations. Although the sample sizes for TDE-H, TDE-He and TDE-featureless classes are small, we find that the timescale distributions of all spectral classes are drawn from the same parent population. van Velzen et al. (2021) had reported that the rise times of their H+He TDEs were longer than the ones for the TDE-H and TDE-He sources. However, this could be due to their smaller sample size along with the small errors from their light curve fitting parameters.
@@
Additionally, we investigated the spectral classes of TDEs in different sub-samples showing either IR emission, repeating flares or X-ray emission. For the IR sample we find that 12 are TDE-H+He, 5 are TDE-H and 1 is TDE-He. For the repeating flare sample we find that 7 are TDE-H+He, 2 are TDE-H and 2 are TDE-featureless. Finally, for the X-ray detection sample we find that 4 are TDE-He, 26 are TDE-H+He, 9 are TDE-H and 3 are TDE-featureless. We estimated the probability that the spectral classes of the different sub-samples are randomly drawn from the full sample using
@@
where is the binomial coefficient. Here, the numerator represents the number of ways to compose a sample with the different spectral classes in each sub-sample while the denominator accounts for all possible selections from the full set of 134 TDEs for the repeating flare sample and 123 TDEs for the X-ray and IR samples 191919126 TDEs have at least X-ray upper limits; however three of them are jetted TDEs with no optical spectrum available; that is why in the denominator for the X-ray calculation. The small p-values in all three cases indicate the spectral composition of these sub-samples is unlikely to be random. This could indicate that the spectral characteristics of TDEs are related to their overall multi-wavelength behaviour.